Today is the 11th anniversary of the terrorist attack that collapsed the World Trade Towers in New York City and killed over 3000 people. Let’s have a moment of silence.
Thank you, now on with our daily report…
- Privacy flags, flags and even more flags
- HYC wants to work with me to find the source of the MDB error. Trying to get schedules synchronized to make that work.
- ldap11 (test replica) exhibited the same failure following Sunday night’s update, so I reverted that one back to the HDB backend yesterday (yeah, I didn’t blog about my work day yesterday, tough!)
- Discovered that when cc is called with both -O2 and -O0 flags, it seems to like the -O2 better (sigh), edit the Makefile, force -O0, rebuild, retest, it still breaks. HYC is now working on a patch.
- Opened ITS#7385 for this issue.
- sharepoint2 and sharepoint3 failed last night. They just up and refused to talk with the backup server. Inquired of the sysadmins if there was a problem.
- Turned in quotes for 300 LTO4 media yesterday. Reduced the usual order by half while we consider the possibility of upgrading the drives to LTO5.
- Still waiting on someone (the lawyers?) to define what has to be vaulted for seven (7) years, so I can adjust the backups and reduce the amount that is going to the vaults – still going to run out of space this year if someone does not make a decision soon.
- SR 49973418 – response from EMC support person (level 1 I assume) that is not relevant and confusing. This should be fun.
- loaded 40 blank tapes into library and labeled them (tested out new script. Yay, it worked and didn’t wipe out all the tapes in the library!)